# Preliminary Investigation into Data Scaling Laws for Imitation Learning-Based End-to-End Autonomous Driving

### v1

@@ Abstract
The end-to-end autonomous driving paradigm has recently attracted lots of attention due to its scalability. However, existing methods are constrained by the limited scale of real-world data, which hinders a comprehensive exploration of the scaling laws associated with end-to-end autonomous driving. To address this issue, we collected substantial data from various driving scenarios and behaviors and conducted an extensive study on the scaling laws of existing imitation learning-based end-to-end autonomous driving paradigms. Specifically, approximately 4 million demonstrations from 23 different scenario types were gathered, amounting to over 30,000 hours of driving demonstrations. We performed open-loop evaluations and closed-loop simulation evaluations in 1,400 diverse driving demonstrations (1,300 for open-loop and 100 for closed-loop) under stringent assessment conditions. Through experimental analysis, we discovered that (1) the performance of the driving model exhibits a power-law relationship with the amount of training data; (2) a small increase in the quantity of long-tailed data can significantly improve the performance for the corresponding scenarios; (3) appropriate scaling of data enables the model to achieve combinatorial generalization in novel scenes and actions. Our results highlight the critical role of data scaling in improving the generalizability of models across diverse autonomous driving scenarios, assuring safe deployment in the real world. Project repository: https://github.com/ucaszyp/Driving-Scaling-Law
@@ 1 Introduction
Some previous works [19, 2] have tried to investigate the effect of scaling up. However, they share the same challenge: insufficient real-world data, resulting in data scaling laws in end-to-end autonomous driving remaining under-explored. As shown in Table. 1, the open-source datasets [4, 5] are typically thousand-scale, far less than the million-scale or billion-scale vision-language data in language models or generative models. Although simulators like CARLA [9] offer a promising solution to the data lack, the huge domain gap hinders its real-world application. Consequently, it remains under-explored whether end-to-end autonomous driving has such common data scaling laws and how can autonomous driving vehicles benefit from the laws.
@@
    Can data scaling endow autonomous driving cars with the generalization to new scenarios and actions?
To answer the questions, we collect and annotate a million-scale dataset named ONE-Drive, which contains over 4 million driving demonstrations (about 30,000 hours) of real-world data in diverse cities and road conditions as shown in Figure. 1. Based on the dataset, we first conduct a detailed analysis of the relationship between training data and model performance. Further, we split the scenarios into 23 types based on the traffic condition and agent behavior to delve into the generalization of the data scaling law. Lastly, we set the training data with different distributions on scenario types to analyze the relationship between data scaling and data distribution. For a comprehensive analysis, we conduct the open-loop test, close-loop test, and real-world deployment to evaluate the planning results.
@@
    Power-laws. Trajectory fitting capability of end-to-end autonomous driving exhibits a power-law relationship with the number of training data (Section 5.1).
@@
    Importance of increasing the quantity of targeted data. Targeted augmentation of long-tailed scenarios exhibits superior efficiency compared to indiscriminate expansion of the overall dataset. (Section 5.2).
@@
    Combinatorial generalization. With the scaling of training data, models gradually acquire combinatorial generalization capabilities, enabling them to combine known information to achieve generalizable planning for new scenarios and actions (Section 5.3).
Table 1: Comparison among the previous dataset and our introduced ONE-Drive for the end-to-end autonomous driving task.
@@
ONE-Drive Real-world  4M
2 Related Work
@@ 2.2 Scaling Laws
3 Method
Figure 3: Distribution of the 23 scenario types.
Following PARA-Drive [33], we implement a parallel modular end-to-end autonomous driving algorithm named ONE model, as shown in Figure. 2. The input is LiDAR point clouds and panoramic RGB images and the expected output is a planning trajectory.
3.1 BEV Encoder
@@
3.2 Parallel Decoder
@@
Ego Motion Planning. In the planning module, to fulfill the requirements of real-world planning for precise adherence to navigation, we employ a multi-layer perceptron (MLP) to encode the traffic light status, road-level route (represented by a collection of dense waypoints at the road level), and the motion state of the ego-vehicle, which are served as contextual information for planning. Following [15, 33], we leverage learnable embeddings to construct ego-vehicle queries. These queries are passed as input to the cross-attention layers to interact with BEV features. Then the queries are used to predict multi-modal future planning trajectories with corresponding scores. During inference, we execute the trajectory with the highest confidence score.
3.3 Loss
The overall loss function is defined as the sum of the task losses:
$\mathcal{L}=\mathcal{L}_{depth}+\mathcal{L}_{map}+\mathcal{L}_{occ}+\mathcal{L}_{static}+\mathcal{L}_{mot}+\mathcal{L}_{plan}$ (1)
where the $L_{depth}$ represents an auxiliary depth loss of forward projection in LSS [27].
4 Experiments Setup
4.1 Implementation Details
@@
Baseline. Our method predicts a 6-second future trajectory with 2-second history information. It uses ResNet34 [10] as the default image backbone. The image resolution is downsampled to $512\times 960$ for training. The default BEV resolution is $232\times 80$ for a perception range of $139.2m\times 48m$ longitudinally and laterally. The default hidden state and embedding dimension are 256. We construct five training datasets with varying numbers of demonstrations: 10 thousand, 50 thousand, 0.7 million, 2 million, and 4 million. We train each model to full convergence during training for a fair comparison. This allows models trained on different dataset sizes to undergo different numbers of training steps. The specific training steps and computational resources are shown in Table. 2
Table 2: Training steps and resources of different training dataset sizes (demonstrations).
@@ 4.2 Data Mining
The distribution of these scenario types in the training dataset (4 million demonstrations) is shown in Figure. 3. For more details about the specific characteristics of the 23 scenario types, please see the supplementary materials.
4.3 Evaluation
@@
The implementation details of the closed-loop simulation, along with the evaluation metrics and visualization results, are presented in the supplementary materials.
5 Experiments
Figure 4: Results of open-loop [(a)] and closed-loop [(b) and (c)] evaluation. (a): Data scaling laws on ADE. Dashed lines represent power-law fits, with the equations and coefficient $r$ provided in the eq. 2. All axes are shown on a logarithmic scale. (b): Scores of each metric in closed-loop evaluation. (c): Total score of closed-loop evaluation. In (b) and (c), the X-axis (number of demonstrations) is displayed on a logarithmic scale and the Y-axis is displayed on a linear scale.
In this section, we investigate the relationship between the scale of training data and the ability of end-to-end driving models to fit expert trajectories (Section 5.1.1). Subsequently, we discover the inconsistency in data scaling laws between open-loop and closed-loop evaluations, revealing that merely increasing data volume does not indefinitely improve autonomous driving planning performance in real-world scenarios (Section 5.1.2). This finding motivates us to explore data scaling strategies beyond simply augmenting data quantity. We examine how data distribution affects the efficiency of data scaling law (Section 5.2). Finally, we demonstrate that end-to-end autonomous driving systems, when scaled appropriately with data, exhibit combinatorial generalization capabilities.
5.1 Unveiling of Data Scaling Laws
5.1.1 Power-law Relationship
Inspired by the power-law data scaling laws observed in large-scale language models [16], we aim to investigate the existence of power-law scaling relationships in imitation learning-based end-to-end autonomous driving. We begin our analysis by examining the Average Displacement Error (ADE) between ground truth trajectories and predicted trajectories in different training dataset scales. Specifically, we utilize open-loop evaluation results from five models with varying training dataset scales, as detailed in Section 3. We designate the amount of training data (quantified by the number of demonstrations used in training) as variable $\rm{X}$ and the normalized trajectory ADE as variable $\rm{Y}$. After applying the logarithmic transformation to both $\rm{X}$ and $\rm{X}$, the presence of a linear relationship would indicate a power-law scaling law in end-to-end autonomous driving data.
As shown in Figure. 4 (a), we conduct the linear model fitting of the log-transformed data. The fitting results yield:
@@ 5.1.1 Power-law Relationship
5.1.2 Differences Between Open-loop And Closed-loop
Given that open-loop evaluation cannot fully reflect a vehicle’s planning performance in real-world scenarios, we further assess the data scaling laws in a closed-loop condition. Specifically, we employ a visual closed-loop simulator to evaluate the vehicle’s closed-loop performance of the models trained on variant scales of data. Following the description in Sec 4.3, we obtain the five closed-loop scores and the total score as illustrated in Figure. 4 (b) and (c).
Through the figure, we observe that nearly all of the closed-loop metrics have a similar relationship with the amount of training data. The score initially increases rapidly with the increase of the data scale, but after reaching an inflection point (about 2 million training demonstrations), the growth rate decelerates as data continues to expand. Surprisingly, this phenomenon differs significantly from those in open-loop assessment.
5.2 Data Quantity and Model Performance
Based on the foundation of data scaling law, we explore the relationship between data quantity and model performance when increasing the scale of training data. Specifically, we categorized our dataset of 2 million demonstrations into 23 scenario types, as defined in Section 4.2, and analyzed the proportion of each type. To simplify the process, we identify two scenario types that exhibit poor performance (DOWN_SUBROAD_LC and SUBROAD_TO_MAINROAD_HIGHWAY) from the 23 categories for in-depth study. By incrementally increasing the representation of these scenario types, we observe the consequent effects on planning performance.
As illustrated in Table. 3, we maintained a constant overall volume of training data while increasing the quantity of these two scenario types in each iteration. Subsequently, we evaluated the model’s open-loop trajectory error after training. The results reveal that for these long-tailed scenarios, doubling the specific training data while keeping the total data volume constant led to improvements in planning performance ranging from 9.7% to 16.9%. Furthermore, quadrupling the specific training data yielded even more substantial gains, with performance improvements between 22.8% and 32.9%. Our findings suggest that even a relatively small increase in scenario-specific data (comprising only several hundred or thousand demonstrations) can lead to significant enhancements in planning performance for these challenging scenarios.
5.3 Combinatorial Generalization
@@
Table 3: Analysis of different scene distribution.
@@
DOWN_SUBROAD_LC 4972 7218 ($+45.1\%$) 21836 ($+339.2\%$) 0.788 0.711 ($-9.7\%$) 0.529 ($-32.9\%$)
SUBROAD_TO_MAINROAD_HIGHWAY 643 1485 ($+130.9\%$) 2569 ($+299.5\%$) 0.786 0.653 ($-16.9\%$) 0.593 ($-22.8\%$)
5.3.1 Quantitative Analysis
To demonstrate the existence of generalization, we split two categories (HIGHWAY_NUDGE_OBS and WAIT_TURN) from the 23 scenario types defined in Sec 4.2 as test categories, while using the remaining 21 categories to form the training set. To be consistent with previous settings, we trained models using three scales of number of training data: 50 thousand, 0.7 million, and 2 million demonstrations. Notably, we applied a strict filtering strategy when selecting test data for these two categories to ensure that each scenario does not overlap with other scenario types. The open-loop test results are shown in Figure. 5. For a convenient comparison, we select scenarios similar to these two scenario types.
We observed that (1) the model trained on 50 thousand demonstrations showed larger displacement error from expert trajectories in the two test scenarios compared to similar scenarios, indicating insufficient generalization capability with small-scale training data. (2) As the training data gradually increased to green demonstrations, the displacement error between the trajectories of the two test scenarios and other scenarios rapidly narrowed. The performance on HIGHWAY_NUDGE_OBS even surpasses the performance on other scenarios that participated in the training. (3) By learning high-speed driving and low-speed nudging obstacles separately from the training data, the model acquired the ability to generalize to high-speed HIGHWAY_NUDGE_OBS scenarios; through learning to turn and queuing at red lights, the model developed the capability to generalize to WAIT_TURN scenarios. Based on these observations, we argue that the data scaling law endows models with the ability to combinatorial generalization.
5.3.2 Qualitative Analysis
To demonstrate the generalization ability more intuitively, we also provide the qualitative visualization of the predicted semantic occupancy in Figure. 6, which shows the planning results of HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios. The BEV maps in the green boxes represent the planning result of the model trained using 50 thousand demonstrations without HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios (50K + Unseen in the legend), the BEV maps in the blue boxes represent the planning result of the model trained using 2 million demonstrations without HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios (2M + Unseen in the legend), and the BEV maps in the orange boxes represent the planning result of the model trained using 2 million demonstrations with HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios (2M + Seen in the legend).

### v2

# Data Scaling Laws for Imitation Learning-Based End-to-End Autonomous Driving

@@ Abstract
The end-to-end autonomous driving paradigm has recently attracted lots of attention due to its scalability. However, existing methods are constrained by the limited scale of real-world data, which hinders a comprehensive exploration of the scaling laws associated with end-to-end autonomous driving. To address this issue, we collected substantial data from various driving scenarios and behaviors and conducted an extensive study on the scaling laws of existing imitation learning-based end-to-end autonomous driving paradigms. Specifically, approximately 4 million demonstrations from 23 different scenario types were gathered, amounting to over 30,000 hours of driving demonstrations. We performed open-loop evaluations and closed-loop simulation evaluations in 1,400 diverse driving demonstrations (1,300 for open-loop and 100 for closed-loop) under stringent assessment conditions. Through experimental analysis, we discovered that (1) the performance of the driving model exhibits a power-law relationship with the amount of data, but this is not the case in closed-loop evaluation. The inconsistency between the two assessments shifts our focus toward the distribution of data rather than merely expanding its volume. (2) a small increase in the quantity of long-tailed data can significantly improve the performance for the corresponding scenarios; (3) appropriate scaling of data enables the model to achieve combinatorial generalization in novel scenes and actions. Our results highlight the critical role of data scaling in improving the generalizability of models across diverse autonomous driving scenarios, assuring safe deployment in the real world.
@@ I Introduction
Some previous works [6, 7] have tried to investigate the effect of scaling up. However, they share the same challenge: insufficient real-world data, resulting in data scaling laws in end-to-end autonomous driving remaining under-explored. As shown in the table. I, the open-source datasets [1, 2] are typically thousand-scale, far less than the million-scale or billion-scale vision-language data in language models or generative models. Although simulators like CARLA [8] offer a promising solution to the data lack, the huge domain gap hinders their real-world application. Consequently, it remains under-explored whether end-to-end autonomous driving has such common data scaling laws and how autonomous driving vehicles can benefit from the laws.
@@
    Can data scaling endow autonomous driving cars with the generalization to new scenarios?
To answer the questions, we collect and annotate a million-scale dataset named ONE-Drive, which contains over 4 million driving demonstrations (about 30,000 hours) of real-world data in diverse cities and road conditions, as shown in the figure. 1. Based on the dataset, we first conduct a detailed analysis of the relationship between training data and model performance. Further, we split the scenarios into 23 types based on the traffic condition and agent behavior to delve into the generalization of the data scaling law. Lastly, we set the training data with different distributions on scenario types to analyze the relationship between data scaling and data distribution. For a comprehensive analysis, we conduct the open-loop test, closed-loop test, and real-world deployment to evaluate the planning results.
@@
    Scaling law. The data scaling law in open-loop evaluation exhibits a power-law relationship with the amount of data, but this is not the case in closed-loop evaluation. The inconsistency between the two evaluations shifts our focus toward the distribution of data rather than merely expanding its volume. (Section V-A).
@@
    Importance of increasing the quantity of targeted data. Targeted augmentation of long-tailed scenarios exhibits superior efficiency compared to indiscriminate expansion of the overall dataset. (Section V-B).
@@
    Combinatorial generalization. With the scaling of training data, models gradually acquire combinatorial generalization capabilities, enabling them to combine known information to achieve generalizable planning for new scenarios and actions (Section V-C).
TABLE I: Comparison between the previous dataset and our introduced ONE-Drive for the end-to-end autonomous driving task.
@@
ONE-Drive (Ours) Real-world  4M
II Related Work
@@ II-B Scaling Laws
III Method
Following PARA-Drive [10], we implement a parallel modular end-to-end autonomous driving algorithm named the ONE model, as shown in the figure. 2. The input is LiDAR point clouds and panoramic RGB images, and the expected output is a planning trajectory.
III-A BEV Encoder
@@
Figure 3: Distribution of the 23 scenario types.
III-B Parallel Decoder
@@
Ego Motion Planning. In the planning module, to fulfill the requirements of real-world planning for precise adherence to navigation, we employ a multi-layer perceptron (MLP) to encode the traffic light status, road-level route (represented by a collection of dense waypoints at the road level), and the motion state of the ego-vehicle, which serve as contextual information for planning. Following [12, 10], we leverage learnable embeddings to construct ego-vehicle queries. These queries are passed as input to the cross-attention layers to interact with BEV features. Then the queries are used to predict multi-modal future planning trajectories with corresponding scores. During inference, we execute the trajectory with the highest confidence score.
Figure 4: Results of open-loop [(a)] and closed-loop [(b) and (c)] evaluation. (a): Data scaling laws on ADE. Dashed lines represent power-law fits, with the equations and coefficient $r$ provided in the eq. 1. All axes are shown on a logarithmic scale. (b): Scores of each metric in closed-loop evaluation. (c): Total score of closed-loop evaluation. In (b) and (c), the X-axis (number of demonstrations) is displayed on a logarithmic scale, and the Y-axis is displayed on a linear scale.
IV Experiments Setup
IV-A Implementation Details
@@
Baseline. Our method predicts a 6-second future trajectory with 2 seconds of historical information. It uses ResNet34 [36] as the default image backbone. The image resolution is downsampled to $512\times 960$ for training. The default BEV resolution is $232\times 80$ for a perception range of $139.2m\times 48m$ longitudinally and laterally. The default hidden state and embedding dimension are 256. We construct five training datasets with varying numbers of demonstrations: 10 thousand, 50 thousand, 0.2 million, 0.7 million, 2 million, and 4 million. We train each model to full convergence during training for a fair comparison. This allows models trained on different dataset sizes to undergo different numbers of training steps. Training steps and computational resources are shown in the table. II
TABLE II: Training steps and resources of different training dataset sizes (demonstrations).
@@ IV-B Data Mining
The distribution of these scenario types in the training dataset (4 million demonstrations) is shown in the figure. 3. For more details about the specific characteristics of the 23 scenario types, please see the supplementary materials.
IV-C Evaluation
@@
V Experiments
In this section, we investigate the relationship between the scale of training data and the ability of end-to-end driving models to fit expert trajectories (Section V-A1). Subsequently, we discover the inconsistency in data scaling laws between open-loop and closed-loop evaluations, revealing that merely increasing data volume does not indefinitely improve autonomous driving planning performance in real-world scenarios (Section V-A2). This finding motivates us to explore data scaling strategies beyond simply augmenting data quantity. We examine how data distribution affects the efficiency of the data scaling law (Section V-B). Finally, we demonstrate that end-to-end autonomous driving systems, when scaled appropriately with data, exhibit combinatorial generalization capabilities.
V-A Unveiling of Data Scaling Laws
V-A1 Power-law Relationship
Inspired by the power-law data scaling laws observed in large-scale language models [3], we aim to investigate the existence of power-law scaling relationships in imitation learning-based end-to-end autonomous driving. We begin our analysis by examining the Average Displacement Error (ADE) between ground truth trajectories and predicted trajectories in different training dataset scales. Specifically, we utilize open-loop evaluation results from five models with varying training dataset scales, as detailed in Section IV. We designate the amount of training data (quantified by the number of demonstrations used in training) as a variable $\rm{X}$ and the normalized trajectory ADE as a variable $\rm{Y}$. After applying the logarithmic transformation to both $\rm{X}$ and $\rm{X}$, the presence of a linear relationship would indicate a power-law scaling law in end-to-end autonomous driving data.
As shown in the figure. 4 (a), we conduct the linear model fitting of the log-transformed data. The fitting results yield:
@@ V-A1 Power-law Relationship
V-A2 Differences Between Open-loop And Closed-loop
Given that open-loop evaluation cannot fully reflect a vehicle’s planning performance in real-world scenarios, we further assess the data scaling laws in a closed-loop condition. Specifically, we employ a visual closed-loop simulator to evaluate the vehicle’s closed-loop performance of the models trained on variant scales of data. Following the description in Sec IV-C, we obtain the five closed-loop scores and the total score as illustrated in the figure. 4 (b) and (c).
From figure (b), we observe distinct relationships between various closed-loop evaluation metrics and the expansion of data volume. Metrics such as navigation and efficiency demonstrate a clear and stable increasing trend, as they are directly related to trajectory and path training. In contrast, the score for the driving rule, while showing an overall upward trend, exhibits instability during its growth. This occurs because driving rules are not explicitly modeled as loss functions in imitation learning—instead, the model must implicitly learn the patterns from large volumes of image-trajectory pairs. Due to inherent limitations in imitation learning, such as causal confusion, performance in rule compliance remains suboptimal. Similarly, metrics including safety and comfort experience rapid improvement initially, but their growth rates gradually slow within the range of 200k to 700k data samples. We hypothesize that in the early stage, increasing data volume enhances the model’s ability to fit trajectories, leading to swift gains in closed-loop safety. In later stages, however, minor discrepancies in trajectory fitting have a diminishing impact on safety in most scenarios. Thus, further improving safety may depend more on optimizing data distribution to enhance performance across diverse situations, rather than simply increasing the amount of data.
V-B Data Quantity and Model Performance
Based on the foundation of the data scaling law, we explore the relationship between data quantity and model performance when increasing the scale of training data. Specifically, we categorized our dataset of 2 million demonstrations into 23 scenario types, as defined in Section IV-B, and analyzed the proportion of each type. To simplify the process, we identify two scenario types that exhibit poor performance (DOWN_SUBROAD_LC and SUBROAD_TO_MAINROAD_HIGHWAY) from the 23 categories for in-depth study. By incrementally increasing the representation of these scenario types, we observe the consequent effects on planning performance.
As illustrated in the table. III, we maintained a constant overall volume of training data while increasing the quantity of these two scenario types in each iteration. Subsequently, we evaluated the model’s open-loop trajectory error after training. The results reveal that for these long-tailed scenarios, doubling the specific training data while keeping the total data volume constant led to improvements in planning performance ranging from 9.7% to 16.9%. Furthermore, quadrupling the specific training data yielded even more substantial gains, with performance improvements between 22.8% and 32.9%. Our findings suggest that even a relatively small increase in scenario-specific data (comprising only several hundred or thousands of demonstrations) can lead to significant enhancements in planning performance for these challenging scenarios.
V-C Combinatorial Generalization
@@
TABLE III: Analysis of different scene distribution.
@@
DOWN_SUBROAD_LC 4972 7218 ($45.1\%$) 21836 ($339.2\%$) 0.788 0.711 ($9.7\%$) 0.529 ($32.9\%$)
SUBROAD_TO_MAINROAD_HIGHWAY 643 1485 ($130.9\%$) 2569 ($299.5\%$) 0.786 0.653 ($16.9\%$) 0.593 ($22.8\%$)
V-C1 Quantitative Analysis
To demonstrate the existence of generalization, we split two categories (HIGHWAY_NUDGE_OBS and WAIT_TURN) from the 23 scenario types defined in Sec IV-B as test categories, while using the remaining 21 categories to form the training set. To be consistent with previous settings, we trained models using three scales of number of training data: 50 thousand, 0.7 million, and 2 million demonstrations. Notably, we applied a strict filtering strategy when selecting test data for these two categories to ensure that each scenario does not overlap with other scenario types. The open-loop test results are shown in the figure. 5. For a convenient comparison, we select scenarios similar to these two scenario types.
We observed that (1) the model trained on 50 thousand demonstrations showed larger displacement error from expert trajectories in the two test scenarios compared to similar scenarios, indicating insufficient generalization capability with small-scale training data. (2) As the training data gradually increased to green demonstrations, the displacement error between the trajectories of the two test scenarios and other scenarios rapidly narrowed. The performance on HIGHWAY_NUDGE_OBS even surpasses the performance on other scenarios that participated in the training. (3) By learning high-speed driving and low-speed nudging obstacles separately from the training data, the model acquired the ability to generalize to high-speed HIGHWAY_NUDGE_OBS scenarios; through learning to turn and queuing at red lights, the model developed the capability to generalize to WAIT_TURN scenarios. Based on these observations, we argue that the data scaling law endows models with the ability to perform combinatorial generalization.
V-C2 Qualitative Analysis
To demonstrate the generalization ability more intuitively, we also provide a qualitative visualization of the predicted trajectory and ground truth map in the figure. 6, which shows the planning results of HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios. The BEV maps in the green boxes represent the planning result of the model trained using 50 thousand demonstrations without HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios (50K + Unseen in the legend), the BEV maps in the blue boxes represent the planning result of the model trained using 2 million demonstrations without HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios (2M + Unseen in the legend), and the BEV maps in the orange boxes represent the planning result of the model trained using 2 million demonstrations with HIGHWAY_NUDGE_OBS and WAIT_TURN scenarios (2M + Seen in the legend).